# Condor: A Code Discriminator Integrating General Semantics with Code Details

### v1

@@ 2.4.3. Dataset Construction
After determining the problem descriptions and the labels corresponding to individual code snippets, we construct data samples in the format $\langle p_{i},c^{neg}_{i},c^{pos}_{i}\rangle$, where $c^{neg}_{i}$ represents an incorrect code submission for problem $p_{i}$, and $c^{pos}_{i}$ represents a correct submission. For a complex problem $p_{i}$, a user often needs multiple attempts to reach the correct solution. For example, with three submissions $S=(s1,s2,s3)$, $s3$ may be the correct code, while $s1$ and $s2$ contain bugs. To ensure that only minimal modifications are required for the model to fix buggy code, we calculate the Jaccard similarity (Jaccard, 1901) between the correct submission $s3$ and the buggy submissions $s1$ and $s2$, denoted as $j(s1,s3)$ and $j(s2,s3)$. We retain only those samples where the Jaccard similarity is greater than 0.9. Given the large number of users, and to ensure diversity in the constructed data samples, we group multiple users into one user group and follow the same process to create data samples. Table 1 shows the basic statistics of the CodeNanoFix dataset. The final dataset consists of training, validation, and test sets, each containing 480, 565, and 120 algorithmic problems and 80,906, 4,742, and 3,583 data samples. The validation set contains algorithmic problems that are generally more specialized, leading to a greater number of problems but fewer data samples. To avoid data leakage, we ensure that problem descriptions in the training and test sets are as distinct as possible, reducing any potential influence of training data on test data. This setup allows for effective evaluation of the model’s performance during training.
2.4.4. Dataset Metric
@@
To verify data samples in the CodeNanoFix dataset requiring only minimal changes to fix bugs, we utilize the relative edit distance (RED) metric to quantify the similarity between buggy code and correct code. We first tokenize the code using the DeepSeek-Coder (Guo et al., 2024) tokenizer and calculate the number of tokens per code sample. Then, we compute the average edit distance (ED) between buggy and correct code versions to quantify the number of modification operations required for repairs. The results indicate that the average edit distance involves significantly fewer tokens than the total token count of the code itself. The RED metric is in the last column of the table, representing the ratio of required edit operations to the original token count (a value between 0 and 1). The specific formula is as follows:
RED=EDAVG(B_Tokens, $\text{RED}=\frac{\text{ED}}{\text{AVG(B\_Tokens, C\_Tokens)}},$
@@
where $AVG$ indicates the average value of two numbers, $B\_Tokens$ represents the number of tokens in the buggy code, and $C\_Tokens$ represents the number of tokens in the correct code. The results show a typical RED of around 0.1, confirming that only minor adjustments are needed to change one code snippet to the other in the CodeNanoFix dataset. As a result, CodeNanoFix is more streamlined and consistent, making it better suitable for evaluating the discriminator’s ability to distinguish subtle changes in code.
3. Experimental Setup
@@ 3.2. Datasets
We conduct experiments on the APPS and MBPP benchmark datasets to further validate Condor’s generalization capabilities. The APPS dataset consists of 5,000 training problems and 5,000 test problems, covering a wide range of complex programming problems. We evaluate each baseline’s original code generation accuracy and its post-process accuracy on the test set, analyzing how Condor performs in more challenging code generation tasks. The MBPP dataset contains 374 training problems, 90 validation problems, and 500 test problems, primarily used for code generation evaluation. Through experiments on the MBPP dataset, we test Condor’s generalization ability and discrimination precision across problems of different scales.
@@
By evaluating Condor on the CodeNanoFix, APPS, and MBPP datasets, we verify the model’s performance across various tasks and data scales, especially its ability to capture code details and generalize to diverse code generation tasks.
3.3. Baselines
We categorize the baselines into two types: baselines for training various discriminators and baselines for code generation. To make the discriminator more accessible, the models involved in training discriminators are typically smaller in parameter size compared to code generation models. Therefore, we select a random strategy, along with models based on CodeRanker, CodeT5, and DeepSeek-Coder-1.3B, as baselines for evaluating the trained discriminator’s performance. To provide a convincing comparison for our discriminator’s performance, we use high-performing models in software engineering tasks, specifically DeepSeek-Coder-Instruct-6.7B, Meta-Llama-3.1-Instruct-8B, and Meta-Llama-3.1-Instruct-70B, as baselines for code generation.
3.3.1. Baselines for Training Discriminators
@@
For the discrimination-related tasks, we select simple strategies and models with smaller parameter sizes as baselines. The random strategy serves as a simple baseline by randomly selecting an output from the model’s generated results as the final output. This approach helps assess the variability of experimental results and ensures the robustness of the benchmark. CodeRanker (Inala et al., 2022) fine-tunes CodeBERT (Feng et al., 2020), an encoder model based on the BERT (Devlin et al., 2018) architecture, for the task of discriminating the outputs of LLMs. Trained with a masked language model on code data, CodeBERT excels in code representation learning, which enhances its applicability to various code understanding and classification tasks. CodeT5 (Wang et al., 2021) is an encoder-decoder structure that incorporates identifier-aware techniques to enhance code comprehension. Following existing studies (Alrashedy et al., 2024; An et al., 2024), we use the encoder of CodeT5 as a baseline for training the discriminator. DeepSeek-Coder-1.3B (Guo et al., 2024) is a decoder model based on the GPT architecture, trained on a large, high-quality code dataset, and performs exceptionally well in various code generation tasks, making it a solid baseline for training discriminators. To ensure that all baselines are compared at a high initial performance level, we generate training data for the MBPP and APPS datasets using code-generation baselines to serve as training samples for the discriminator. This approach not only provides high-quality training examples for the discriminator but also ensures consistency in the training data across all discriminators, facilitating a fair comparison.
3.3.2. Baselines for Code Generation
@@
For code generation and repair tasks, we select high-performing, large-scale GPT-style models as baselines. DeepSeek-Coder-Instruct-6.7B (Guo et al., 2024) offers stronger code understanding and generation capabilities than its 1.3B version, making it a high-performance baseline for generation. Meta-Llama-3.1-Instruct-8B (Dubey et al., 2024), with 8 billion parameters, is a powerful code generation model based on the LLaMA architecture (Touvron et al., 2023; Roziere et al., 2023), specifically optimized for complex code generation scenarios. Meta-Llama-3.1-Instruct-70B, with 70 billion parameters, handles even greater code generation complexity, establishing itself as a top-performing baseline for code generation tasks.
3.4. Metrics
@@
To comprehensively evaluate the performance of the Condor model in code classification and discrimination tasks, we use multiple evaluation metrics, including Precision, Recall, F1-score, and Pass@1 (Powers, 2020) to measure the model’s performance before and after discriminating.
3.4.1. Precision
@@ 3.4.3. F1-score
[…]
@@ 3.4.4. Pass@1
3.5. Implementation Details
We train two versions of Condor with different parameter scales (i.e., Condor-110M and Condor-1.3B) based on these strategies. We choose to use CodeT5’s encoder and DeepSeek-Coder-1.3B as the base models to train Condor at different parameter scales for two main reasons. First, both models demonstrate strong performance in various software engineering tasks and are widely used in the field, providing a solid foundation. Second, compared to larger models like DeepSeekCoder-Instruct-6.7B or Code-Llama-3.1-Instruct-70B, they are relatively small, which helps us control the parameter size of the discriminator, making training and deployment more efficient across a variety of tasks. The margin value $m$ for the contrastive learning loss function is set to 2 to reflect the typical magnitude of semantic differences between correct and incorrect code.
We conduct our experiments on a machine running the Ubuntu operating system, equipped with 8 NVIDIA A100 GPUs, each with 80GB of memory, and 500GB of system memory. This hardware setup provides the necessary support for large-scale model training and dataset loading. To ensure the reproducibility of our experiments, we implement the models and experiments using the Transformers and Huggingface (Huggingface, 2024b) libraries. Additionally, to accelerate the training process, we utilize DeepSpeed (Rajbhandari et al., 2020) and Accelerate (Huggingface, 2024a) libraries, which optimize computational efficiency in a multi-GPU environment, significantly reducing training time.
4. Results
4.1. RQ1: Effectiveness of Condor
@@
4.1.1. Effectiveness on Classification Scenario
Table 2. Condor’s Performance in Classification.
Model Precision Recall F1-Score
CodeRanker (110M) 59.42 77.56 67.29
CodeT5 (110M) 60.09 76.95 67.48
DeepSeek-Coder (1.3B) 71.99 62.77 67.06
Condor (110M) 65.99 80.04 72.34
Condor (1.3B) 74.39 72.40 73.38
Table 2 shows the classification performance of the models with different parameter sizes on the CodeNanoFix dataset. From the table, it can be observed that smaller models (such as CodeRanker and CodeT5) have lower precision but higher recall, while larger models (such as DeepSeek-Code) demonstrate higher precision but lower recall, achieving more balanced overall performance.
Two factors can explain these results. First, smaller models, with their limited numbers of parameters, struggle to distinguish subtle code differences accurately. As a result, they tend to broadly label potentially correct code to capture as many functionally correct samples as possible, leading to higher recall. However, this broad labeling strategy also increases the likelihood of false positives, resulting in lower precision. For example, CodeRanker and CodeT5 have recall values of 77.56% and 76.95%, respectively, but their precision only reaches 59.42% and 60.09%. Second, larger models, with more parameters, can capture finer-grained code features, making them more precise in identifying correct code and thus improving precision. However, as these models become more conservative, they are more cautious when labeling code as correct, potentially missing some functionally correct but borderline samples, leading to lower recall. For example, DeepSeek-Coder (1.3B) achieves a precision of 71.99% but a recall of only 62.77%.
Condor introduces embedding-level and data-level strategies to capture the code details, enhancing the model’s sensitivity to functional differences in code, and significantly improving the overall performance of models of various sizes. These strategies enable Condor to better capture subtle code changes, improving precision and recall for both small and large models. Specifically, Condor (110M) boosts precision to 65.99% and recall to 80.04%, showing a substantial improvement over CodeRanker and CodeT5. Meanwhile, Condor (1.3B) increases precision to 74.39% and recall to 72.40%, achieving a more balanced performance, with an F1-Score of 73.38%, higher than all other models. Overall, regardless of model size, Condor’s code-detail-aware strategies achieve significant improvements in both precision and recall, resulting in enhanced overall performance.
4.1.2. Effectiveness on Discrimination Scenario
Table 3. Condor’s discriminative performance in the CodeNanoFix dataset.

### v2

@@ II-D3 Dataset Construction
After determining the problem descriptions and the labels corresponding to individual code snippets, we construct data samples in the format $\langle p_{i},c^{neg}_{i},c^{pos}_{i}\rangle$, where $c^{neg}_{i}$ represents an incorrect code submission for problem $p_{i}$, and $c^{pos}_{i}$ represents a correct submission. For a complex problem $p_{i}$, a user often needs multiple attempts to reach the correct solution. For example, with three submissions $S=(s1,s2,s3)$, $s3$ may be the correct code, while $s1$ and $s2$ contain bugs. To ensure that only minimal modifications are required for the model to fix buggy code, we calculate the Jaccard similarity [jaccard1901] between the correct submission $s3$ and the buggy submissions $s1$ and $s2$, denoted as $j(s1,s3)$ and $j(s2,s3)$. We retain only those samples where the Jaccard similarity is greater than 0.9. Given the large number of users, and to ensure diversity in the constructed data samples, we group multiple users into one user group and follow the same process to create data samples. Table I shows the basic statistics of the CodeNanoFix dataset. The final dataset consists of training, validation, and test sets, each containing 480, 565, and 120 algorithmic problems and 80,906, 4,742, and 3,583 data samples. The validation set contains algorithmic problems that are generally more specialized, leading to a greater number of problems but fewer data samples. To avoid data leakage, we ensure that problem descriptions in the training and test sets are as distinct as possible, reducing any potential influence of training data on test data. This setup allows for effective evaluation of the model’s performance during training.
II-D4 Dataset Metric
@@
To verify data samples in the CodeNanoFix dataset requiring only minimal changes to fix bugs, we utilize the relative edit distance (RED) metric to quantify the similarity between buggy code and correct code. We first tokenize the code using the DeepSeek (DS)-Coder [guo2024deepseekcoder] tokenizer and calculate the number of tokens per code sample. Then, we compute the average edit distance (ED) between buggy and correct code versions to quantify the number of modification operations required for repairs. The results indicate that the average edit distance involves significantly fewer tokens than the total token count of the code itself. The RED metric is in the last column of the table, representing the ratio of required edit operations to the original token count (a value between 0 and 1). The specific formula is as follows:
$\text{RED}=\frac{\text{ED}}{\text{AVG(B\_Tokens, C\_Tokens)}},$
@@
where $AVG$ indicates the average value of two numbers, B_TokensB\_Tokens represents the number of tokens in the buggy code, and C_TokensC\_Tokens represents the number of tokens in the correct code. The results show a typical RED of around 0.1, confirming that only minor adjustments are needed to change one code snippet to the other in the CodeNanoFix dataset. As a result, CodeNanoFix is more streamlined and consistent, making it better suitable for evaluating the discriminator’s ability to distinguish subtle changes in code.
III Experimental Setup
@@ III-B Datasets
We conduct experiments on the APPS [apps], MBPP [mbpp], and LiveCodeBench [jain2024livecodebench] benchmark datasets to further validate Condor’s generalization capabilities. The APPS dataset consists of 5,000 training problems and 5,000 test problems, covering a wide range of complex programming problems. We evaluate each baseline’s original code generation accuracy and its post-process accuracy on the test set, analyzing how Condor performs in more challenging code generation tasks. The MBPP dataset contains 374 training problems, 90 validation problems, and 500 test problems, primarily used for code generation evaluation. Through experiments on the MBPP dataset, we test Condor’s generalization ability and discrimination precision across problems of different scales. In addition, we evaluate on LiveCodeBench (version 5), which comprises 880 problems focused on dynamic and real-world coding scenarios. Unlike APPS and MBPP, LiveCodeBench does not provide a dedicated training set for discriminators, making it particularly suitable for testing the generalization ability of Condor and all baselines in practical, unconstrained settings.
@@
By evaluating Condor on the CodeNanoFix, APPS, MBPP, and LiveCodeBench datasets, we verify the model’s performance across various tasks and data scales, especially its ability to capture code details and generalize to diverse code generation tasks.
III-C Baselines
We categorize the baselines into two groups: those for evaluating discriminators and those for evaluating code generation performance. For the discriminator baselines, we select smaller-scale models for supervised discriminator training [codebert2020, codet52021, coderanker], and utilize larger-scale models for few-shot discrimination. This design ensures that the trained discriminators remain lightweight, while also allowing us to assess the few-shot reasoning abilities of more powerful models. Specifically, we include a random selection strategy, as well as models based on CodeRanker, CodeT5, and DS-Coder (1.3B, where ‘DS’ represents ‘DeepSeek’) as baselines to assess the trained discriminator’s performance. For few-shot discrimination, we additionally consider DS-Coder-Instruct (1.3B) and Llama-3.1-Instruct (8B). In contrast, to provide a robust upper bound and a compelling comparison for our discriminator’s effectiveness, we adopt high-performing large models in software engineering tasks as code generation baselines, including DS-Coder-Instruct (6.7B), Llama-3.1-Instruct (8B), and Llama-3.1-Instruct (70B).
III-C1 Baselines for Training Discriminators
@@
For discrimination-related tasks, we adopt three main baseline settings to comprehensively evaluate model performance: (1) a heuristic-based random selection strategy, (2) discriminators trained using small-scale models, and (3) few-shot discrimination using instruction-following LLMs. The random strategy provides a simple baseline by randomly selecting an output from the model’s generated candidates, offering a lower bound for benchmark performance. For supervised discrimination, we focus on lightweight models, such as CodeRanker (built on CodeBERT), the encoder-only variant of CodeT5, and DS-Coder-1.3B. These models are selected for their proven efficiency and ease of deployment, aligning with real-world scenarios where a smaller discriminator re-ranks outputs from a larger generator. This design is particularly effective in resource-constrained environments, where deploying a large generation model alone is impractical [coderanker, codet52021, guo2024deepseekcoder]. CodeRanker [coderanker] fine-tunes CodeBERT [codebert2020], an encoder model based on the BERT [bert2018] architecture, for the task of discriminating the outputs of LLMs. Trained with a masked language model on code data, CodeBERT excels in code representation learning, which enhances its applicability to various code understanding and classification tasks. CodeT5 [codet52021] is an encoder-decoder structure that incorporates identifier-aware techniques to enhance code comprehension [divot5, twinxsql]. Following existing studies [alrashedy2024codet5_encoder, an2024codet5_encoder], we use the encoder of CodeT5 as a baseline for training the discriminator. DS-Coder-1.3B [guo2024deepseekcoder] is a decoder model based on the GPT architecture, trained on a large, high-quality code dataset, and performs exceptionally well in various code generation tasks, making it a solid baseline for training discriminators [grammarcoder].
To ensure all trained baselines are compared under equally strong initial conditions, we augment the training set beyond CodeNanoFix by incorporating training data from the MBPP and APPS datasets. For these datasets, we use our code generation baselines to produce candidate outputs, which then serve as high-quality training samples for the discriminator. This strategy not only enriches the training data with diverse and challenging examples but also guarantees consistency across all discriminator baselines, thereby facilitating a fair and robust comparison.
In addition, we incorporate a few-shot discrimination setting using instruction-following LLMs, specifically DS-Coder-Instruct (1.3B) and Llama-3.1-Instruct (8B), where the model is prompted to judge code correctness or select the correct answer from multiple candidates [wang2024advanced]. The temperature is set to 0 to enforce deterministic predictions. Two common prompt formats are used: (1) “You are given a question and a candidate answer. Your task is to determine whether the answer correctly solves or responds to the question. EXAMPLES. TASK.” and (2) “You are given a question and a list of generated outputs. Your task is to select the most likely correct answer from the list. EXAMPLES. TASK.” The EXAMPLES and TASK denote the few-shot examples provided and the specific inference task given a particular input, respectively. This design covers both practical deployment scenarios and advanced LLM settings, enabling a thorough and fair comparison of different discrimination approaches.
III-C2 Baselines for Code Generation
@@
For code generation and repair tasks, we select high-performing, large-scale GPT-style models as baselines. DS-Coder-Instruc (6.7B) [guo2024deepseekcoder] offers stronger code understanding and generation capabilities than its 1.3B version, making it a high-performance baseline for generation. Llama-3.1-Instruc (8B) [dubey2024llama3], with 8 billion parameters, is a powerful code generation model based on the LLaMA architecture [touvron2023llama, roziere2023codellama], specifically optimized for complex code generation scenarios. Llama-3.1-Instruct (70B), with 70 billion parameters, handles even greater code generation complexity, establishing itself as a top-performing baseline for code generation tasks. For each input question, every model generates 10 candidate answers. In the subsequent discrimination stage, the discriminator must select the single most likely correct output from these 10 candidates as the final answer. This setup enables a direct evaluation of the discrimination model’s ability to identify the best solution among strong generative outputs.
III-D Metrics
@@
To comprehensively evaluate the performance of the Condor model in code classification and discrimination tasks, we use multiple evaluation metrics, including Precision, Recall, F1-score, and Pass@1 [powers2020evaluationprecisionrecallfmeasure] to measure the model’s performance before and after discriminating.
III-D1 Precision
@@ III-D3 F1-score
$\text{F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@ III-D4 Pass@1
III-E Implementation Details
We train two versions of Condor with different parameter scales (i.e., Condor-110M and Condor-1.3B) based on these strategies. We choose to use CodeT5’s encoder and DS-Coder (1.3B) as the base models to train Condor at different parameter scales for two main reasons. First, both models demonstrate strong performance in various software engineering tasks and are widely used in the field, providing a solid foundation. Second, compared to larger models like DS-Coder-Instruct (6.7B) or Llama-3.1-Instruct (70B), they are relatively small, which helps us control the parameter size of the discriminator, making training and deployment more efficient across a variety of tasks. The margin value $m$ for the contrastive learning loss function is set to 2 to reflect the typical magnitude of semantic differences between correct and incorrect code.
We conduct our experiments on a machine running the Ubuntu operating system, equipped with 8 NVIDIA A100 GPUs, each with 80GB of memory, and 500GB of system memory. This hardware setup provides the necessary support for large-scale model training and dataset loading. To ensure the reproducibility of our experiments, we implement the models and experiments using the Transformers and Huggingface [huggingface] libraries. Additionally, to accelerate the training process, we utilize DeepSpeed [deepspeed] and Accelerate [acclerate] libraries, which optimize computational efficiency in a multi-GPU environment, significantly reducing training time.
IV Results
[…]
IV-A RQ1: Effectiveness of Condor
@@
IV-A1 Effectiveness on Classification Scenario
TABLE II: Condor’s Performance in Classification and Discrimination Scenarios, where ‘FS’ indicates a few-shot manner and ‘MP’ represents the max potential. The ‘w/o’ indicates without a specific strategy in the ablation study process.
Model Classification Discrimination
Precision Recall F1-Score
Pass@1 of DS-Coder
-Instruct (6.7B, MP: 75.41)
Pass@1 of Llama-3.1
-Instruct (8B, MP: 68.66)
Pass@1 of Llama-3.1
-Instruct (70B, MP: 83.70)
Original 50.00 50.00 50.00 43.23 36.81 52.64
Random 50.18 50.93 50.55 43.23 36.75 53.39
CodeRanker (110M) 59.42 77.56 67.29 42.84 37.48 57.69
CodeT5 (110M) 60.09 76.95 67.48 42.19 37.65 56.93
DS-Coder (1.3B) 71.99 62.77 67.06 43.54 37.32 58.90
DS-Coder-Instruct (1.3B, FS) 50.88 25.71 34.16 43.18 36.70 53.34
Llama-3.1-Instruct (8B, FS) 55.45 90.26 68.70 42.59 38.88 55.74
Condor (110M) 65.99 80.04 72.34 43.73 39.60 59.64
Condor (1.3B) 74.39 72.40α 73.38 50.85 45.27 62.63
w/o Contrastive Learning 73.12 72.06α 72.59 48.03 42.37 60.95
w/o Intermediate Data 67.98 74.66 71.16 45.29 38.54 59.84
The left side of the Table II presents the classification performance of various approaches on the CodeNanoFix dataset, including our Condor model, random selection, models trained specifically for the discrimination task, and large language models used in a few-shot setting without dedicated training. Across all evaluated metrics, Condor consistently outperforms both the random baseline and the supervised discriminators, as well as few-shot LLMs.
For models trained specifically for discrimination, small models such as CodeRanker and CodeT5 (both 110M parameters) achieve recall rates of about 77% but their precision is only around 60%, resulting in F1-scores just above 67%. This indicates that small models, due to their limited parameter capacity, tend to broadly classify code as correct in order to maximize recall. While this helps capture more true positives, it also leads to a higher rate of false positives, reducing overall precision. In contrast, larger models like DS-Coder (1.3B) benefit from increased representational power, achieving a higher precision of about 72% and a recall of 63%, with a similar F1-score. The performance of larger models is relatively more balanced, as they can more precisely identify correct code while being more cautious and filtering out borderline cases.
Few-shot LLMs, such as Llama-3.1-Instruct (8B, FS) and DS-Coder-Instruct (1.3B, FS), show a different pattern. Without task-specific training, these models tend to favor high recall—Llama-3.1-Instruct (8B, FS) achieves a recall of over 90%, but their precision drops to around 55%, resulting in an F1-score below 69%. This suggests that few-shot prompting helps these models identify most of the correct cases but also causes them to misclassify many incorrect samples as correct, leading to more false positives. DS-Coder-Instruct (1.3B, FS) performs even worse, with both precision and recall dropping below 51% and 26%, respectively, highlighting the limitations of relying on general generation capabilities without targeted discrimination training.
In contrast, Condor incorporates embedding-level and data-level strategies specifically designed to enhance the model’s sensitivity to code semantics and structural details. As a result, Condor (110M) substantially improves precision to 66%, an improvement of about 6 percentage points compared to traditional discriminators like CodeRanker (59%) and CodeT5 (60%). At the 1.3B scale, Condor reaches 74% precision and 72% recall, with an F1-score of 73%, surpassing DS-Coder (1.3B) by more than 9 percentage points in the recall metric. These improvements are primarily due to Condor’s detail-aware representations, which allow the model to more accurately identify subtle distinctions between correct and incorrect code. This enables Condor to maintain high recall while greatly reducing false positives, thus boosting precision.
IV-A2 Effectiveness on Discrimination Scenario
The right side of Table II reports the discrimination results, measuring the functional correctness of the top-1 code selected by each discriminator using the Pass@1 metric.
We first analyze whether subtle code differences impact model performance. Specifically, we examine label (i.e., correct or error) consistency in Llama-3.1-Instruct (70B) when addressing the same input with similar tokens (Jaccard similarity bigger than 0.7). The results show that 47.55% of such samples have inconsistent labels (i.e., the probability of making mistakes). Separately, we investigate the Jaccard similarity for cases with label inconsistencies and find that over 37.44% of such samples have a Jaccard similarity exceeding 0.7. This shows that even a 70B-parameter model can generate incorrect code due to subtle differences, underscoring the need for discriminators to capture fine-grained code details.
After applying discriminators to the generated outputs, we observe clear performance improvements. Notably, Condor (1.3B) consistently achieves the largest gains across all code generation models. For DS-Coder-Instruct (6.7B), Pass@1 rises from 43.23% to 50.85%, an increase of 17.6% relative to the baseline. On Llama-3.1-Instruct (70B), Condor (1.3B) boosts Pass@1 from 52.64% to 62.63%, corresponding to a 19% relative improvement. Condor (110M) also yields improvements across all models, though the magnitude is smaller, which reflects the importance of sufficient model capacity for capturing nuanced code features in complex generation settings. By comparison, other discriminators such as CodeRanker and CodeT5 may offer some improvement in Pass@1, but the gains are not consistent across all settings. The random selection baseline performs nearly identically to the original models, underscoring that naive selection strategies contribute little to overall performance.
Answer to RQ1: The experimental results demonstrate that code details play a crucial role in improving both classification and discrimination performance. By introducing code detail-aware strategies, Condor significantly enhances its ability to capture subtle code differences, leading to improved precision and recall across different models. For example, Condor (1.3B) achieves a precision of 74.39% and a recall of 72.40%. In the discrimination scenario, Condor’s performance is particularly outstanding. On Llama-3.1 Instruct 70B, Condor (1.3B) increases Pass@1 from 52.64% to 62.63%. This demonstrates Condor’s ability to identify correct code and exhibit strong discriminative capabilities across models of various sizes.
IV-B RQ2: Generalizability to Code Generation Tasks
To comprehensively evaluate the generalization capability of Condor, we conduct experiments on three representative datasets: APPS, MBPP, and the recently introduced LiveCodeBench (version 5). Table III, Table IV, and Table V present the Pass@1 results of various discriminators, including both traditional baselines and our Condor model.
TABLE III: Condor’s generalization capability in APPS, where ‘FS’ indicates a few-shot manner and ‘MP’ represents the max potential.